### v1

@@ II-B Intersection Point
The IP is the point in space where the net GRF vectors at adjacent time steps intersect. Fig. 1 illustrates the IP concept for quiet stance and kneeling. Unlike the COP or COM, the use of the IP takes into account the translational and angular accelerations of the body and the IP height is used to quantify the neuromuscular control strategy.
The IP height, denoted by $h_{\text{IP}}$, is calculated using the COP and GRF data obtained from force plate measurements. The angle between the GRF vector and the vertical direction is denoted as $q(t)$ at time $t$; see Fig. 1. Assuming subtle movements of the body and using a small angle approximation, we obtain
@@
We will further discuss how to process $h_{\text{IP}}(t)$ in frequency domain to analyze the balance performance in Section IV.
@@ II-C Neural Balance Control
Following observations from [11], we obtain the physiological meaning of three sets of parameters (i.e., $\alpha_{i}$, ${\boldsymbol{\beta}}_{i}$, ${\boldsymbol{\sigma}}_{i}$) that are associated with the LQR-based neural controller: (i) $\alpha_{i}$ characterizes the control effort employed to attain stability, (ii) ${\boldsymbol{\beta}}_{i}$ penalizes the contribution of ankle, knee, and hip joints control, and (iii) ${\boldsymbol{\sigma}}_{i}$ signifies the neuromuscular impedance and noise at each joint. The output of (7) is considered as $y_{1i}=-F^{x}_{i}$, $i=t,d$, and $y_{2t}=\tau_{1}$ (ankle torque for stance) or $y_{2d}=\tau_{2}$ (knee torque for kneeling) that are directly measured from force plate. The negative sign in $y_{1i}$ is due to the convention that a positive horizontal GRF is directed towards the positive $x$-axis; see Fig. 1. From (3) and using the Jacobian matrix, we obtain
@@ III-A Experiment Setup and Protocol
Fig. 2 shows the experimental setup for quiet stance and kneeling gaits in construction. A wearable inertial measurement unit (IMU) (from LP-RESEARCH Inc.) was attached to the subject’s chest to measure postural sway accelerations. A force plate (from Bertec Corporation) was used to collect GRF and COP data. An extra pulse sensor was used to capture the subject’s heart rate during experiments. The human subjects were also equipped with a bilateral knee exoskeleton with quasi-direct drive actuators. The exoskeleton possessed high-torque, high-backdrivability, and high-bandwidth features [29]. The design was compact and light-weight (unilateral unit weighed only $1.7$ kg, including electronics and battery) with high-torque output ($35$ Nm peak torque), large range of motion (00-$160$ deg flexion), and high rotation speed ($16.2$ rad/s). All sensor data were synchronized and collected at $100$ Hz on a portable high-performance micro-processor (Intel NUC7i7DNK, Intel Corp.).
A VR system (HTC Vive Pro) was employed to immerse the subjects in diverse visual environments during test trials. A virtual construction site was built with the Unity game engine to simulate the typical low-elevation (LE) and high-elevation (HE) construction work environments. Fig. 3 shows the VR viewpoint for the quiet stance and Fig. 3 shows the VR scene for welding task. The LE scenario depicted the subject standing on a small platform slightly elevated from the ground. The HE scenario on the other hand simulated the subject standing atop a $30$-story building under ongoing construction.
@@ III-A1 Protocol A - Posture balance evaluation
Subjects were asked to perform a quiet stance and kneeling task while maintaining balance in a simulated environment for $60$ s. The visual stimuli varied between trials by altering the LE and HE. This setup facilitated two distinct trials, each designed to assess the impact of elevation perception on balance. The wearable knee exoskeleton was introduced in a subsequent variation.
The research work in [30] has shown that the knee joint’s deviation from a natural stance is small and a stiffness-based controller effectively reduces muscle fatigue and enhances balance. Consequently, a stiffness knee torque controller $\tau_{e}=-k_{r}\theta_{2}$ is used for the quiet stance gait, where $\tau_{e}$ is the exoskeleton output torque, $k_{r}$ represents the reference joint stiffness, and $\theta_{2}$ denotes the knee joint’s deviation from a natural stance. In contrast, during kneeling, the knee joint acts as a pin joint, bearing a significant portion of the body’s weight. This increased loading on the knee joint leads to large variations in joint angles and requires great effort to maintain balance. Inspired by the work in [31], the inertia and Coriolis terms in (1) are neglected and a proportional-derivative (PD) knee assistive torque controller with gravity compensation is considered
@@ III-A2 Protocol B - Welding task assessment
A welding trade was considered as welders often have to hold awkward gaits while conducting tasks with precision requirements. The dual task of maintaining upright balance and performing welding accurately is challenging. Protocol B introduced a welding task in the VR scene, see Fig. 3, simulating a dynamic and task-focused assessment. As shown in Fig. 2, participants were equipped with a VR controller in their dominant hand, emulating a welding gun, and were tasked with welding a workpiece as accurately as possible within the VR environment. To enhance realism, welding gloves were provided, and no time constraints were imposed. The welding task was visually represented by ink emitted from the weld gun tip upon contact with the workpiece, resulting in a visual weld line. The completed works were documented as screenshots for subsequent analysis.
Similar to Protocol A, subjects underwent quiet stance and kneeling with four TOIs in Protocol B: (i) TOI 5: Welding task in an LE without the exoskeleton; (ii) TOI 6: Welding task in an HE without the exoskeleton; (iii) TOI 7: Welding task in an LE with the exoskeleton; and (iv) TOI 8: Welding task in an HE with the exoskeleton. Participants were given test trials to get familiar with the welding setup and operation. Participants were allowed a $10$-min rest period between each trial to prevent task-induced fatigue and ensure accurate task execution. In both Protocols A and B, the orders of the trials were randomized for evaluation purposes.
@@ III-B Data Collection and Processing Methods
During the experiments, the $x$- and $y$-axis were respectively defined as anterior-posterior (A-P), medial-lateral (M-L) directions. Fig. 2 illustrates two key measurements used in analysis to quantify movement variability and postural stability: the $95\%$ acceleration ellipse area $A_{i}^{\text{Acc}}$ and the $95\%$ sway ellipse area $A_{i}^{\text{Sway}}$, $i=t,d$. $A_{i}^{\text{Acc}}$ is calculated as the elliptical area that contains 95% of the acceleration variations (from IMU) in the A-P ($\ddot{x}^{m}_{i}$) and M-L ($\ddot{y}^{m}_{i}$) planes. It quantifies the variability in body accelerations and thus reflects the dynamic control and execution of movement. In contrast, $A_{i}^{\text{Sway}}$ is obtained as the elliptical area that contains $95\%$ COP data (from force plate) in the A-P ($x_{\text{CP}}$) and M-L ($y_{\text{CP}}$) directions. It represents the spatial variability of the COP movements, offering insights into the stability of postural equilibrium and the subject’s ability to maintain or regain balance.
@@
Subjective perceptions were also collected by pre- and post-experiment questionnaires (provided in the supplemental materials). Pre-experimentally, subjects’ inherent fear responses were quantified using the James Geer Fear (JGF) questionnaire [32], and the Cohen Acrophobia (CA) questionnaire [33], aimed at understanding their reactions to heights. Both questionnaires used a five-point Likert scale [34]. A subsequent post-experiment questionnaire sought to elucidate subjects’ experiences within the VR scene and their perceptions of the experimental procedures, aiming to refine and enhance the experimental design through iterative feedback. This dual approach, melding subjective feedback with objective data, provides a holistic understanding of the impact of experimental conditions on both the psychological and physiological dimensions of participant response.
@@ IV Data Analysis and Simulation
Fig. 6 shows an example of the IP frequency curve normalized by $h_{\text{COM}}=z^{m}_{i}$. We focus on the crossover frequency (CF) of the IP height frequency curve with the COM height (i.e., ratio $h_{\text{IP}}/h_{\text{COM}}$ at one), denoted as $\omega^{\text{CF}}$, and the slope of the high-frequency asymptote (HFA), denoted as $k^{\text{HFA}}$. To determine and calculate $k^{\text{HFA}}$ from the IP curves, an exponential fit was applied. For kneeling, the IP generally lies above the COM and it is challenging to define the CF and HFA in the same way as in quiet stance. To maintain consistency in the analysis of balance control strategies, we introduce the concept of virtual CF and HFA by normalizing the IP height as a fraction of the upright stance COM height; see Fig. 1.
@@
The IP frequency curves offer insights into human balance control strategies. At low frequencies, the IP height greater than the COM height indicates active neuromuscular engagement and high control effort to maintain balance. As the frequency increases, the IP curve decays, suggesting a transition from active to passive control strategies, characterized by $\omega^{\text{CF}}$. The HFA slope $k^{\text{HFA}}$ of the IP curve represents the passive control mechanism dominant at higher frequencies, reflecting the neuromechanical impedance of the joints.
@@
We also conducted simulations to analyze the impact of various neural balance controller parameter sets on the IP height frequency curve. These simulations provided a controlled environment to investigate the complex interplay between various biomechanical and neurological factors contributing to human balance control. The simulation allowed for the exploration of a wide range of scenarios and conditions that may be difficult or impossible to test experimentally. To simulate and analyze the impact of control parameters on human balance strategies, the TIP and DIP models with the LQR-based neural controller with initial condition ${\boldsymbol{x}}(0)={\boldsymbol{0}}$ to compute IP height frequency behaviors. Table I lists the parameter values for the models that were determined by typical human biomechanical data [35]. The simulation involved executing a series of trials to analyze the model’s response under various conditions, each lasting $50$ s with data sampled at a frequency of $100$ Hz. To ensure robust statistical analysis, $30$ trials were conducted for each parameter set.
@@ V Results
V-A Experimental Results
@@
Figs. 7-7 show the IP frequency curves for TOIs 1 to 4. The IP height magnitude is above the COM at lower frequencies, demonstrating greater control effort at these frequencies. The IP curve decays with increasing frequency, indicating a reduced influence of active control. The IP curve also reaches an asymptote below the COM at high frequencies, suggesting a transition to passive control mechanisms. The spikes in the IP frequency curve indicate changes in balance strategy, which is direct effort of an individual attempting to stabilize the gait. On examining each individual’s IP curve across TOIs, several observations are obtained. In the balance tests, for instance, subject S3 generated high frequency spikes in the IP curves when provided with exoskeleton assistance. The observation suggests a counter-intuitive perception of instability, causing the subjects to alter their balance strategies to maintain balance. Subject S6 also showed increased variation in the IP curve with the exoskeleton. On the other hand, presence of the exoskeleton reduced both high-frequency and high-magnitude spikes for S4. For S1, there were large spikes at for LE cases at low frequencies, while smaller spikes were observed for HE cases at high frequencies, indicating a shift in balance strategy on experiencing altered visual stimuli. It is evident that the interplay between HE and exoskeleton use did not uniformly impact all individuals; some subjects such as S5 and S7 maintained a consistent and stable balance strategy that remained unchanged across all balance TOIs, regardless of elevation or exoskeleton.
Figs. 7-7 show the IP curves generated by subjects performing quiet stance while conducting welding tasks. The near absence of high frequency spikes is noted. Examination of these curves across individual subjects reveals that focusing on tasks (TOIs 5 to 8) generally enhances postural stability. When compared to the quiet balance, most subjects in kneeling gait exhibited smoother asymptotic behavior in the IP frequency response, which suggests a stable control strategy employed during the welding tasks. This improved stability can be a result of the subjects focusing on the welding task, permitting a stable balance strategy to be instinctively chosen. An exception to this trend is observed in the case of subject S1, whose IP curve does not follow the typical characteristics expected of healthy human quiet stance. This can be attributed to the subject modifying their gait during task performance, transitioning from upright stance to a quarter-squat position. Thus, IP curves have the potential to discriminate between different types of gait, underscoring the advantages of exploring IP behavior in other common construction worker gaits.
Fig. 8 shows the IP height-frequency curves for the kneeling balance task and the kneeling welding task for 6 subjects. All kneeling curves are similar to the cases of quiet stance. The relative IP height in relation to the COM is overall higher in kneeling than in stance. The kneeling IP curves give insight into individual balance strategies in response to different testing stimuli. For the balance experiments, HE increased the magnitude of high frequency spikes in S8 and S11 and reduced the spikes in S6. The exoskeleton assistance reduced and increased high frequency spikes in LE balance tests for S8 and S9, respectively. When outfitted with the exoskeleton, S6, S7 and S10 show smoother IP curves during welding task performance. S9’s balance strategy varied greatly across all TOIs with high frequency spikes in all cases except TOI 4. Overall, S6, S8 and S11 have smoother IP curves for TOIs 6 to 8 than those in TOIs 1 to 4.
@@
Table II lists the mean values of $\omega_{i}^{\text{CF}}$ and $k_{i}^{\text{HFA}}$ through one-way ANOVA for all TOIs. Upon the introduction of the exoskeleton, there was a discernible decrease in the mean CF $\bar{\omega}_{t}^{\mathrm{CF}}$ during the balance experiments, suggesting a shift in the balance control strategy employed by the subjects. The value of $\bar{\omega}_{t}^{\mathrm{CF}}$ was also found to increase with the HE visuals. For the welding task scenario, overall $\bar{\omega}_{t}^{\mathrm{CF}}$ is lower than that of the balance scenario, indicating the introduction of the task reduced balance control effort. While the combination of the exoskeleton and HE visuals lowered $\bar{\omega}_{t}^{\text{CF}}$ in the balancing experiments, performing welding tasks increased its value. The mean slope of the HFA $|\bar{k}_{t}^{\text{HFA}}|$ had a higher magnitude for the balance experiments than the welding task experiments. The condition of HE without exoskeleton had the lowest magnitude of $|\bar{k}_{t}^{\text{HFA}}|$, while and LE with exoskeleton support had the highest $|\bar{k}_{t}^{\text{HFA}}|$ in balancing experiments. The vice-versa was true for the welding task performance experiments. The kneeling IP descriptors are also presented in Table. II. The value of $\bar{\omega}^{\text{CF}}_{d}$ does not vary with visual stimuli. It however increases with the exoskeleton assistance and welding task. The task performance increases the magnitude of $|\bar{k}_{d}^{\text{HFA}}|$ for HE without exoskeleton assistance and LE with exoskeleton assistance. For LE without exoskeleton support and HE with exoskeleton support, the values of $|\bar{k}_{d}^{\text{HFA}}|$ however decreased.
Fig. 9 shows the PSD curves under various test conditions. The figures compare the PSD results across the balance and welding task during stance and kneeling gaits. By examining the PSD plots, we can concisely analyze the postural sway behavior. For example, low postural sway is observed in welding task TOIs (asymptote begins around $0.4$-$0.6$ Hz) compared to the balance TOIs (asymptote begins around $3$-$4$ Hz), regardless of the gait performed. Stance overall has higher postural sway than kneeling gait as showcased by the PSD frequency curves located at higher frequencies. High-elevation visual feedback caused more postural sway during quiet stance than kneeling gait as seen by the higher area under the curve. While the exoskeleton reduced the sway in most TOIs, it allows for increased sway in TOIs 7 and 8 during kneeling. As shown in Figs. 7 and 8, the IP methodology works to complement these observations by providing more detailed physiological and joint-level insights, as detailed in Section V-B.
@@
Fig. 10 shows an example of assistive torque $\tau_{e}$ by the exoskeleton during a 10-s interval for both stance and kneeling gaits in TOI 4 from one subject. The stiffness-based controller during quiet stance generated extension torques and the gravity-compensating PD controller for kneeling demonstrated dynamic torque range. These torque outputs helped maintain the knee joint’s desired position and stability. The effectiveness of the exoskeleton assistance in improving balance can be further observed through metrics such as the $A_{d}^{\text{Sway}}$ and $A_{d}^{\text{Acc}}$, as shown in Fig. 11. The reduced ellipse areas in the presence of the exoskeleton, particularly during HE conditions and welding tasks, indicate that the device contributes to better postural stability and reduced body sway.

### v2

@@ II-B Intersection Point
The IP is the point in space where the net GRF vectors at adjacent time steps intersect. Fig. 1 illustrates the IP concept for quiet stance and kneeling. The IP height, denoted by $h_{\text{IP}}$, is calculated using the COP and GRF data obtained from force plate measurements. The angle between the GRF vector and the vertical direction is denoted as $q(t)$ at time $t$; see Fig. 1. Assuming subtle movements of the body and using a small angle approximation, we obtain
@@
Unlike traditional methods such as COP or COM analysis, the IP model accounts for both translational and rotational accelerations of the body, providing a comprehensive perspective on balance control strategies. By incorporating both aspects of body motion, IP analysis offers a holistic understanding of how individuals maintain postural stability. Specifically, IP height across different frequencies serves as a crucial metric for quantifying neuromuscular control efforts [4]. The frequency analysis is essential because balance control involves movements at different speeds. Low-frequency components reflect active, slower, and large-amplitude adjustments like swaying to maintain posture, while high-frequency elements capture passive, rapid, and small-amplitude corrections in response to quick perturbations [29]. Analyzing these different frequency components helps understand how the body’s control mechanisms shift.
IP height is directly related to this frequency analysis because it changes systematically across different frequency bands [4]. At low frequencies, the IP height is typically above the COM, indicating that the control strategy emphasizes active rotational stability through the hip and ankle joints. This corresponds to slower, larger adjustments in posture, where the body consciously engages in active neuromuscular control, such as muscle co-contractions and sensory feedback, to counterbalance significant deviations from an upright stance. As the frequency increases, the IP height moves below the COM. This shift reflects a change in control strategy toward passive, rapid, fine-tuned adjustments, primarily through the lower body. At these higher frequencies, the body relies less on active control and more on the passive mechanical properties of the muscles and joints, such as joint stiffness and elasticity, to manage small, quick perturbations and maintain stability.
@@ II-C Neural Balance Control
Following observations from [11], the physiological effects of balance control strategies at the joint level can be inferred from the LQR-based neural controller controller parameters $\alpha_{i}$, ${\boldsymbol{\beta}}_{i}$, ${\boldsymbol{\sigma}}_{i}$. Specifically, $\alpha_{i}$ characterizes the control effort employed to attain stability, ${\boldsymbol{\beta}}_{i}$ represents the contribution of ankle, knee, and hip joints to balance control, and ${\boldsymbol{\sigma}}_{i}$ captures the neuromuscular impedance and noise at each joint.
The output of (7) is considered as $y_{1i}=-F_{x}^{i}$, $i=t,d$, and $y_{2t}=\tau_{1}$ (ankle torque for stance) or $y_{2d}=\tau_{2}$ (knee torque for kneeling) that are directly measured from force plate. The negative sign in $y_{1i}$ is due to the convention that a positive horizontal GRF is directed towards the positive $x$-axis; see Fig. 1. From (3) and using the Jacobian matrix, we obtain
@@ III-A Experiment Setup and Protocol
Fig. 2 shows the experimental setup for quiet stance and kneeling gaits in construction. A wearable inertial measurement unit (IMU) (from LP-RESEARCH Inc.) was attached to the subject’s chest to measure postural sway accelerations. A force plate (from Bertec Corporation) was used to collect GRF and COP data. A heart rate sensor (from World Famous Electronics LLC.) was used to capture the subject’s heart rate during experiments. The human subjects were also equipped with bilateral knee exoskeletons to provide assistive torques and distribute musculoskeletal load and stress from the knee joint to the thigh and shank. The exoskeleton’s weight is supported by a waist belt, which is connected to the thigh support frame through elastic straps. The belt also contains an integrated compartment for the microcontroller and battery. The elastic straps secure the knee actuator, ensuring proper alignment. The thigh support frame includes a height-adjustable aluminum linkage on the lateral side of the leg, along with cuffs positioned on the back of the upper thigh and front of the lower thigh. The knee joint actuation system possessed high-torque, high-backdrivability, and high-bandwidth features [30].
The portable knee exoskeleton features a compact and lightweight design (unilateral unit weighed only $1.7$ kg, including electronics and battery); see Fig. 2. Each leg has a quasi-direct drive actuator ($35$ Nm peak torque) to provide a large range of motion (00-$160$ deg flexion), and high rotation speed ($16.2$ rad/s). The exoskeleton torque control was implemented on a low-level microcontroller (from PJRC Teensy) via CAN bus. All sensor data were synchronized and collected at 100 Hz using custom Python scripts running on a portable high-performance microcomputer (model NUC7i7DNK, Intel Corp.) The system ensures precise time-stamping of sensor data, enabling accurate synchronization during post-processing and analysis.
A VR system (HTC Vive Pro) was employed to immerse the subjects in diverse visual environments during test trials. A virtual construction site was built with the Unity game engine to simulate the typical low-elevation (LE) and high-elevation (HE) construction work environments. Fig. 3 shows the VR viewpoint for the quiet stance and Fig. 3 shows the VR scene for welding tasks. The LE scenario depicted the subject standing on a small platform slightly elevated from the ground. The HE scenario on the other hand simulated the subject standing atop a $30$-story building under ongoing construction.
@@ III-A1 Protocol A - Posture balance evaluation
Subjects were asked to perform a quiet stance and kneeling task while maintaining balance in a simulated environment for $60$ s. Subjects were given motion training to mimic construction worker gaits based on consultations with construction professionals. No additional contact points such as hands or other body parts were allowed during the experiment tasks. The visual stimuli varied between trials by altering the LE and HE. This setup facilitated two distinct trials, each designed to assess the impact of elevation perception on balance. The wearable knee exoskeleton was introduced in a subsequent variation.
The results in [31] have shown that the knee joint’s deviation from a natural stance is small and a stiffness-based controller effectively reduces muscle fatigue and enhances balance. Consequently, a stiffness knee torque controller $\tau_{e}=-k_{r}\theta_{2}$ was used for the quiet stance, where $\tau_{e}$ is the exoskeleton output torque, $k_{r}$ represents the reference joint stiffness, and $\theta_{2}$ denotes the knee joint’s deviation from a natural stance. In contrast, during kneeling, the knee joint acts as a pin joint, bearing a significant portion of the body’s weight. This increased loading on the knee joint leads to large variations in joint angles and requires great effort to maintain balance. Inspired by the work in [32], the inertia and Coriolis terms in (1) are neglected and a proportional-derivative (PD) knee assistive torque controller with gravity compensation is considered
@@ III-A2 Protocol B - Welding task assessment
A welding trade was considered as welders often have to hold awkward gaits while conducting tasks with precision requirements. The dual task of maintaining upright balance and performing welding accurately is challenging. Protocol B introduced a welding task in the VR scene (see Fig. 3) and generated a dynamic and task-focused assessment. As shown in Fig. 2, participants were equipped with a VR controller in their dominant hand to emulate a welding gun, and were asked to weld a workpiece as accurately as possible within the VR environment. To enhance realism, welding gloves were provided and no time constraints were imposed. The welding task was visually represented by ink emitted from the weld gun tip upon contact with the workpiece, resulting in a visual weld line. The completed works were documented as screenshots for subsequent analysis.
Similar to Protocol A, subjects underwent quiet stance and kneeling with four TOIs in Protocol B: (i) TOI 5: Welding task in an LE without the exoskeleton; (ii) TOI 6: Welding task in an HE without the exoskeleton; (iii) TOI 7: Welding task in an LE with the exoskeleton; and (iv) TOI 8: Welding task in an HE with the exoskeleton. Participants were given test trials to familiarize themselves with the welding setup and operation. They were instructed to perform balance and welding tasks exclusively in quiet stance and kneeling gaits, without adopting any other postures. Participants were allowed a $10$-min rest period between each trial to prevent task-induced fatigue and ensure accurate task execution. In both Protocols A and B, the orders of the trials were randomized for evaluation purposes.
In this study, participants wore the exoskeleton for approximately $60$ min across various tasks and experimental conditions. Although individual trials were relatively short, with each lasting around $60$ s, the cumulative exposure time allowed for the assessment of the device’s comfort and usability beyond initial wear effects. Participants were provided with rest breaks between trials to minimize fatigue. To evaluate safety and comfort, both quantitative measures (e.g., postural sway, task performance, and physiological data such as heart rate) and qualitative feedback were gathered through post-experiment questionnaires, where subjects provided input on their comfort levels and overall experience with the exoskeleton.
@@ III-B Data Collection and Processing Methods
During the experiments, the $x$- and $y$-axis were respectively defined as anterior-posterior (A-P) and medial-lateral (M-L) directions. Fig. 2 illustrates two key measurements used in analysis to quantify movement variability and postural stability: the $95\%$ acceleration ellipse area $A_{i}^{\text{Acc}}$ and the $95\%$ sway ellipse area $A_{i}^{\text{Sway}}$, $i=t,d$. $A_{i}^{\text{Acc}}$ is calculated as the elliptical area that contains 95% of the acceleration variations (from IMU) in the A-P ($\ddot{x}^{m}_{i}$) and M-L ($\ddot{y}^{m}_{i}$) planes. It quantifies the variability in body accelerations and thus reflects the dynamic control and execution of movement. In contrast, $A_{i}^{\text{Sway}}$ is obtained as the elliptical area that contains $95\%$ COP data (from force plate) in the A-P ($x_{\text{CP}}$) and M-L ($y_{\text{CP}}$) directions. It represents the spatial variability of the COP movements, offering insights into the stability of postural equilibrium and the subject’s ability to maintain or regain balance.
@@
Subjective perceptions were also collected by pre- and post-experiment questionnaires (provided in the supplemental materials). Pre-experimentally, subjects’ inherent fear responses were quantified using the James Geer Fear (JGF) questionnaire [33], and the Cohen Acrophobia (CA) questionnaire [34], aimed at understanding their reactions to heights. Both questionnaires used a five-point Likert scale [35]. A subsequent post-experiment questionnaire sought to elucidate subjects’ experiences within the VR scene, their perceptions of the experimental procedures, and their comfort and experience with the exoskeleton. This comprehensive questionnaire approach aimed to refine and enhance the experimental design through iterative feedback, while also gathering specific insights on the usability of the wearable device. This dual approach, melding subjective feedback with objective data, provides a holistic understanding of the impact of experimental conditions on both the psychological and physiological dimensions of participant response, as well as the practicality and user acceptance of the exoskeleton technology in the experimental context.
@@ IV Data Analysis and Simulation
Fig. 6 shows an example of the IP frequency curve normalized by $h_{\text{COM}}=z^{m}_{i}$. We focus on the crossover frequency (CF) of the IP height frequency curve with the COM height (i.e., ratio $h_{\text{IP}}/h_{\text{COM}}$ at one), denoted by $\omega^{\text{CF}}$, and the slope of the high-frequency asymptote (HFA), denoted by $k^{\text{HFA}}$. To determine and calculate $k^{\text{HFA}}$ from the IP curves, an exponential fit was applied. For kneeling, the IP generally lies above the COM and it is challenging to define the CF and HFA in the same way as in stance. To maintain consistency in the analysis of balance control strategies, we introduce the concept of virtual CF and HFA by normalizing the IP height as a fraction of the upright stance COM height; see Fig. 1.
@@
The IP frequency curves offer insights into human balance control strategies. At low frequencies, the IP height greater than the COM height indicates active neuromuscular engagement and high control effort to maintain balance. As the frequency increases, the IP curve decays, suggesting a transition from active to passive control strategies, characterized by $\omega^{\text{CF}}$. The HFA slope $k^{\text{HFA}}$ of the IP curve represents the passive control mechanism dominant at higher frequencies, reflecting the neuromechanical impedance of the joints. This variation in IP height across frequencies provides insights into how muscles and joints adapt to different balance demands—offering information beyond what traditional COP or COM analysis can capture. By examining IP height at different frequencies, it becomes possible to understand how individuals adjust their balance strategies in response to various conditions, such as task complexity or elevation changes.
@@
We also conducted simulations to analyze the impact of various neural balance controller parameter sets on the IP height frequency curve. The simulation allowed for the exploration of a wide range of scenarios and conditions that may be difficult or impossible to test experimentally. To simulate and analyze the impact of control parameters on human balance strategies, the TIP and DIP models with the LQR-based neural controller with zero initial condition to compute IP height frequency behaviors. Table I lists the parameter values for the models that are determined by typical human biomechanical data [36]. The simulation involved executing a series of trials to analyze the model’s response under various conditions, each lasting $50$ s with data sampled at a frequency of $100$ Hz. To ensure robust statistical analysis, $30$ trials were conducted for each parameter set.
@@ V Results
Figs. 7-7 show the IP frequency curves for TOIs 1 to 4. The IP height magnitude is above the COM at low frequencies, demonstrating highly active control effort at these frequencies. The IP curve decays with increasing frequency, indicating a reduced influence of active control. The IP curve also reaches an asymptote below the COM at high frequencies, suggesting a transition to passive control mechanisms. The spikes in the IP frequency curve indicate changes in balance strategy, which is the direct effort of an individual attempting to stabilize the gait. On examining each individual’s IP curve across TOIs, several observations are obtained. In the balance tests, for instance, subject S3 generated high-frequency spikes in the IP curves with exoskeleton assistance. The observation suggests a counter-intuitive perception of instability, causing the subjects to alter their balance strategies to maintain balance. Subject S6 also showed increased variation in the IP curve with the exoskeleton. On the other hand, the presence of the exoskeleton reduced both high-frequency and high-magnitude spikes for S4. For S1, large spikes are observed for LE cases at low frequencies, while small spikes were observed for HE cases at high frequencies, indicating a shift in balance strategy on experiencing altered visual stimuli. It is evident that the interplay between HE and exoskeleton use did not uniformly impact all individuals; some subjects (e.g., S5 and S7) maintained a consistent and stable balance strategy across all balance TOIs, regardless of elevation or exoskeleton usage.
Figs. 7-7 show the IP curves generated by subjects performing quiet stance while conducting welding tasks. Examination of these curves across individual subjects reveals that focusing on tasks (TOIs 5 to 8) generally enhances postural stability. When compared to the quiet balance, most subjects in kneeling gait exhibited smoother asymptotic behavior in the IP frequency response, which suggests a stable control strategy employed during the welding tasks. This improved stability can be a result of the subjects focusing on the welding task, permitting a stable balance strategy to be instinctively chosen. An exception to this trend is observed in the case of subject S1, whose IP curve does not follow the typical characteristics expected of healthy human quiet stance. This can be attributed to the subject modifying their gait during task performance, transitioning from an upright stance to a quarter-squat position. Thus, IP curves have the potential to discriminate between different types of gait.
@@
Fig. 8 shows the IP height-frequency curves for the kneeling balance task and the kneeling welding task for 6 subjects. All kneeling curves are similar to the cases of quiet stance. The IP height in relation to the COM is overall higher in kneeling than in stance. The kneeling IP curves give insight into individual balance strategies in response to different testing stimuli. For the balance experiments, HE increased the magnitude of high-frequency spikes in S8 and S11 and reduced the spikes in S6. The exoskeleton assistance reduced and increased high-frequency spikes in LE balance tests for S8 and S9, respectively. When outfitted with the exoskeleton, S6, S7, and S10 show smoother IP curves during welding task. S9’s balance strategy varied greatly across all TOIs with high-frequency spikes in all cases except TOI 4. Overall, S6, S8, and S11 have smoother IP curves for TOIs 6 to 8 than those in TOIs 1 to 4.
@@
The one-way analysis of variance (ANOVA) was conducted to analyze the differences in $\omega_{i}^{\text{CF}}$ and $k_{i}^{\text{HFA}}$ across all TOIs. Table II presents the resulting mean values and statistical significance of these parameters. Upon the introduction of the exoskeleton, there was a discernible decrease in the mean CF $\bar{\omega}_{t}^{\mathrm{CF}}$ during the balance experiments. This suggests a shift in the balance control strategy employed by the subjects. The value of $\bar{\omega}_{t}^{\mathrm{CF}}$ was also found to increase with the HE visuals. For the welding task scenario, overall $\bar{\omega}_{t}^{\mathrm{CF}}$ is lower than that of the balance scenario, indicating the introduction of the task reduced balance control effort. While the combination of the exoskeleton and HE visuals lowered $\bar{\omega}_{t}^{\text{CF}}$ in the balancing experiments, performing welding tasks increased its value. The mean slope of the HFA $|\bar{k}_{t}^{\text{HFA}}|$ had a higher magnitude for the balance experiments than the welding task experiments. The condition of HE without exoskeleton had the lowest magnitude of $|\bar{k}_{t}^{\text{HFA}}|$, while LE with exoskeleton support had the highest $|\bar{k}_{t}^{\text{HFA}}|$ in balancing experiments. The vice-versa was true for the welding task performance experiments. The kneeling IP descriptors are also presented in Table. II. The value of $\bar{\omega}^{\text{CF}}_{d}$ does not vary with visual stimuli. It however increases with the exoskeleton assistance and welding task. The task performance increases the magnitude of $|\bar{k}_{d}^{\text{HFA}}|$ for HE without exoskeleton assistance and LE with exoskeleton assistance. For LE without exoskeleton support and HE with exoskeleton support, the values of $|\bar{k}_{d}^{\text{HFA}}|$ however decreased.
Fig. 9 shows the PSD curves under various test conditions. The figures compare the PSD results across the balance and welding tasks during stance and kneeling gaits. By examining the PSD plots, we can concisely analyze the postural sway behavior. For example, low postural sway is observed in welding task TOIs (asymptote begins around $0.4$-$0.6$ Hz) compared to the balance TOIs (asymptote begins around $3$-$4$ Hz), regardless of the gait performed. Stance overall has higher postural sway than kneeling gait as showcased by the PSD frequency curves located at high frequencies. High-elevation visual feedback caused more postural sway during stance than kneeling gait as seen by the higher area under the curve. While the exoskeleton reduced the sway in most TOIs, it allowed for increased sway in TOIs 7 and 8 during kneeling. As shown in Figs. 7 and 8, the IP metric complements these observations by providing detailed physiological and joint-level insights.
@@
Fig. 10 shows an example of assistive torque $\tau_{e}$ by the exoskeleton during a 10-s interval for both stance and kneeling gaits in TOI 4 from one subject. The stiffness-based controller during quiet stance generated extension torques and the PD controller for kneeling demonstrated dynamic torque range. These torque outputs helped maintain the knee joint’s desired position and stability. The effectiveness of the exoskeleton assistance in improving balance can be further observed through metrics such as the $A_{d}^{\text{Sway}}$ and $A_{d}^{\text{Acc}}$, as shown in Fig. 11. The reduced ellipse areas in the presence of the exoskeleton, particularly during HE conditions and welding tasks, indicate that the device contributes to better postural stability and reduced body sway.